Stuart Weddie's father is 55 years old and wants to set up a cash flow stream that would be forever. He would like to receive $15,000 every year, beginning at the end of this year. If he could invest in account earning 9 percent, how much would he have to invest today to receive his perpetual cash flow? (Round to the nearest dollar.) 

$200,000 

$222,222 

$135.200 

$166,667 
Answer #1

Amount required each year = P = $15000

Required Rate of Return = r = 9%

Let the amount invested today be PV

Hence, PV = P/(1+r) + P/(1+r)2 + ..... = P/r = 15000/0.09 = $166,667